OpenStudy (anonymous):

Solve each equation. 4 = x2 -7

OpenStudy (john_es):

Do you mean \[4=x^2-7\]?

OpenStudy (anonymous):

yes sir !

OpenStudy (john_es):

Group numbers in one side, \[4+7=x^2\Rightarrow 11=x^2\] Then, take the square root of both members, \[\pm\sqrt{11}=x\]